# UEFA Champions League entry is incentive incompatible again from the 2024/25 season

### v4

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification for the post-2024 UEFA Champions League is not always based on sporting merit. Consequently, a match with misaligned incentives has been barely avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended to safeguard the integrity of the competition. Fairness can be easily guaranteed by reversing the order of two articles in the Champions League regulation. Since UEFA repeatedly commits the same mistake that is widely known among researchers, governing bodies in sports are called to work more closely together with the scientific community.
@@
Hoc est, quod ait Heraclitus: “In idem flumen bis descendimus et non descendimus.111 Source: https://www.perseus.tufts.edu/hopper/text?doc=Perseus%3Atext%3A2007.01.0080%3Aletter%3D58
@@ 1 Introduction
Ensuring the fairness and integrity of competitions is a key responsibility of tournament organisers (Kendall and Lenten,, 2017; Devriesere et al.,, 2024). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League in 2024 with the following words (UEFA, 2024a, ):
@@
However, is the UEFA Champions League qualification based on sporting merit under any circumstances? Is there any issue in its rules that threatens the fairness of this tournament and can potentially jeopardise competition?
The current note uncovers that, while UEFA may be committed to sporting merit in principles, their regulations do not avoid an unfair game in a domestic league. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots are assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig.
The problem is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, its solution has been known since March 2013 when the first version of Dagaev and Sonin, (2018) became available.
@@ 5 Conclusions
The issue presented above provides an instructive lesson for sports governing bodies around the world: commitment to sporting merit requires not only media releases but appropriately designed competitions. Since rule changes such as the introduction of the European performance spot can always affect fairness, a thorough assessment is necessary before they are implemented. In order to achieve this fundamental aim, Kendall and Lenten, (2017) suggest two alternative calls to action:
@@
Hopefully, UEFA and other sports governing bodies will adopt one of these options in the future. This would be especially important because, as it has been highlighted in Section 2, UEFA repeatedly commits the same mistake in allocating vacant slots—while the fair solution has been provided by Dagaev and Sonin, (2018), and the first version of this paper is available since March 2013.

### v5

# Misaligned incentives in the post-2024 UEFA Champions League: A managerial failure

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification for the post-2024 UEFA Champions League is not always based on sporting merit as we show: a match with misaligned incentives has barely been avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended to safeguard the integrity of the competition. Fairness can be easily guaranteed by reversing the order of two articles in the Champions League regulation. UEFA repeatedly commits the same mistake in allocating vacant slots, which seems to indicate a clear managerial failure, especially because the solution is widely known. Thus, governing bodies in sports are called to work more closely together with the scientific community.
@@
Hoc est, quod ait Heraclitus: “In idem flumen bis descendimus et non descendimus.”111 Source: https://www.perseus.tufts.edu/hopper/text?doc=Perseus%3Atext%3A2007.01.0080%3Aletter%3D58
@@ 1 Introduction
Ensuring the fairness and integrity of sports is one of the key responsibilities of tournament organisers (Kendall and Lenten,, 2017; Devriesere et al.,, 2024). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League in 2024 with the following words (UEFA, 2024a, ):
@@
However, is the UEFA Champions League qualification based on sporting merit under any circumstances? Is there any issue in its rules that threatens the fairness of this prestigious tournament and can potentially jeopardise competition?
The current paper uncovers that, while UEFA may be committed to sporting merit in principles, their regulations do not avoid an unfair game in a domestic league. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots are assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig.
The problem is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, the root of the problem resides in the allocation of vacant slots, which led to some instances of incentive incompatibility in the past, even though its solution has been known since March 2013 when the first version of Dagaev and Sonin, (2018) became available. Making the same mistake repeatedly seems to indicate a clear managerial failure.
@@ 5 Conclusions
Table 2: Timeline of incentive incompatibility in UEFA club competitions
caused by the allocation of vacant slots
Date/Period Issue
2009/10–2014/15 Incentive incompatibility of the UEFA Europa League entry (Dagaev and Sonin,, 2018)
May 2012 The match Feyenoord vs Heerenveen is played, where Heerenveen was better off by losing than by playing a draw (Csató,, 2021, Chapter 2.1)
March 2013 The first version of Dagaev and Sonin, (2018) becomes available that verifies the incentive incompatibility of the UEFA Europa League entry and offers a general solution to the problem
2015/16–2017/18 Incentive incompatibility of the UEFA Champions League entry (Csató,, 2019)
2015/16–2023/24 Incentive incompatibility of the UEFA Champions League seeding (Csató, 2020b, )
2024/25– Incentive incompatibility of the UEFA Champions League entry (this paper)
The problem identified is especially worrying because it is rooted in the allocation of vacant slots, a topic that was repeatedly responsible for incentive incompatibility in UEFA club competitions as summarised in Table 2—while the fair solution has already been provided by Dagaev and Sonin, (2018). This issue offers an instructive lesson for decision-makers around the world: commitment to sporting merit requires not only media releases but appropriately designed tournaments. Since rule changes such as the introduction of the European performance spot can always affect fairness, a thorough assessment is necessary before they are implemented. In order to achieve this fundamental aim, Kendall and Lenten, (2017) suggest two alternative calls to action:
@@
Hopefully, most sports governing bodies, including UEFA, will adopt one of these options in the near future.